OpenStudy (anonymous):

what are the coordinates of the center of the circle shown below? x^2+y^2 -8x -10y-8

OpenStudy (anonymous):

4,5

OpenStudy (lgbasallote):

group all the x and y firs \[(x^2 -8x) + (y^2 - 10y) = 8\] now complete the square of each group @NotSObright mind backing up your answer? /:)

OpenStudy (anonymous):

(x^2-8x+16)+(y^2-10y+100)-16-100-8=0 (x-4)^2+(y-5)^2=124

OpenStudy (anonymous):

thank you guys